An urn contains 8 balls numbered 1-8 and a ball is randomly drawn from the urn. Let

A = “Ball number 2, 4 or 6 is chosen” and let B = “Ball number 3,4,5,6 or 7 is

chosen”. Calculate the following probabilities:

a)P(A)

b) P(A')

c) P(A∪B)

d) P(A|B)
Answer #1

a)P(A)=3/8 (as there are 3 favorable outcome out of 8)

b)P(A')=1-P(A)=1-3/8=5/8

c)P(A u B)=6/8=3/4 (as there are 6 favorable outcome out of 8)

d)P(A|B) =2/5 (as there are 2 outcomes 4,6 are among 5 outcome of B)